# Strategic Planning and Rationalizing on Trees Make LLMs Better Debaters

### v1

@@ Abstract
Winning competitive debates requires sophisticated reasoning and argument skills. There are unique challenges in the competitive debate: (1) The time constraints force debaters to make strategic choices about which points to pursue rather than covering all possible arguments; (2) The persuasiveness of the debate relies on the back-and-forth interaction between arguments, which a single final game status cannot evaluate. To address these challenges, we propose TreeDebater, a novel debate framework that excels in competitive debate. We introduce two tree structures: the Rehearsal Tree and Debate Flow Tree. The Rehearsal Tree anticipates the attack and defenses to evaluate the strength of the claim, while the Debate Flow Tree tracks the debate status to identify the active actions. TreeDebater allocates its time budget among candidate actions and uses the speech time controller and feedback from the simulated audience to revise its statement. The human evaluation on both the stage-level and the debate-level comparison shows that our TreeDebater outperforms the state-of-the-art multi-agent debate system. Further investigation shows that TreeDebater shows better strategies in limiting time to important debate actions, aligning with the strategies of human debate experts.
@@ 1 Introduction
Competitive debate is a structured battleground of arguments. It plays a crucial role in domains such as education (Huryn, 1986; Ball, 2021), legislation (O’Connell, 2002), and politics (Hart & Jarvis, 1997; Holbrook, 1999). In a debate competition, two sides argued on the same motion under strict time limits. It requires complex reasoning skills and precise selection of impactful arguments. Expert debaters allocate time wisely among arguments in different stages. Unlike other tasks, there is no ground-truth answer in a debate: The winner is determined by who sways the audience or judges more effectively. Research shows that success in debates depends heavily on a debater’s ability of intense back-and-forth competition and on-the-fly decision-making: predict, respond to, and adjust based on their opponent’s arguments (Rapoport, 1960; William L. Benoit & Verser, 2003; Salvi et al., 2024). Although large language models (LLMs) are effective in generating arguments, human evaluators continue to rate AI debaters as less convincing than human opponents in live debates (Flamino et al., 2025). These gaps highlight that current LLMs struggle with planning effective arguments in the dynamic back-and-forth interaction of competitive debate.
@@
The main challenges of debate AI lie in strict time limits and the lack of objective reward signals. First, the timed setting makes it impossible for LLMs to generate verbose arguments for each candidate action. For example, there can be a trade-off between attacking the opponent’s claims and defending one’s own claims. A formal competitive debate typically consists of three stages: an opening statement, a rebuttal statement, and a closing statement. In each stage, the debater is required to carefully allocate their limited speaking time to propose new claims, reinforce the existing claims, attack the opponent’s claims, or rebut the opponent’s attack. The debaters must prioritize the most impactful actions and dynamically adjust their plan as the debate progresses. However, unlike games such as Go (Silver et al., 2016, 2017) or Werewolf (Xu et al., 2023, 2024), there are no rule-based winning conditions. Instead, the winner depends on the evolving flow of arguments and counterarguments between the two sides. This dynamic nature makes it difficult to plan based on the final reward signals.
@@
In this work, we propose to model the dynamic interaction on trees and show that planning on these trees can significantly improve the debating capacity of LLMs. Human debate experts implicitly employ tree-like reasoning: they rehearse potential rebuttals before the debate starts by anticipating the claims an opponent might say and formulating the potential responses for each of them in a tree format. During the debate, they track the flow of the debate with another tree to keep a structured mental map of which points have been addressed or remain standing. Inspired by these human strategies, we introduce a new debate system TreeDebater to help LLMs make tactical decisions in the debate. Akin to human debaters, we introduce two kinds of trees in TreeDebater: Rehearsal Tree to hypothesize the opponent’s attacks in advance, and Debate Flow Tree to track the debate status. Specifically, in the rehearsal trees, TreeDebater scores arguments based on the potential tree-structured attacks and defenses. During the debate, TreeDebater keeps track of the debate status through the debate flow trees and provide the candidate action set. TreeDebater retrieves prepared arguments from the rehearsal tree for each candidate action to draft the statement. Finally, TreeDebater revises its statement based on simulated audience feedback and uses a speech-based search algorithm to fit the statement into the speaking time limitation.
@@
To evaluate the effectiveness of TreeDebater in competitive debate, we carried out extensive experiments with the multi-agent debate system Agent4Debate (Zhang et al., 2024), which has shown comparable results to human debaters. We design two kinds of human evaluation to compare the debate performance: stage-level head-to-head comparison with a fixed debate context, and debate-level end-to-end comparison with an Oxford-style debate format. Human evaluation demonstrates that TreeDebater consistently outperforms the baseline in the dimension of the average persuasiveness score and the stage-level and debate-level win rate. It is preferred 1.5x and 3.5x over the baseline in the stage-level and debate-level comparison with the Gemini-2.0-flash backbone, where the gain also generalizes to another backbone, DeepSeek-V3.
@@ 2 Related Work
Computational Argumentation. Previous studies used computational methods to study argumentation processes, such as argument mining (Lawrence & Reed, 2019), the polarity of arguments (Agarwal et al., 2022; Liu et al., 2021), argument structure (Zhang et al., 2016; Li et al., 2020), and argument generation (Hua et al., 2019; Lin et al., 2024). Slonim et al. (2021) built the first autonomous debating system, Project Debater, with four argumentation modules: argument mining, argument knowledge base, argument rebuttal, and debate construction. While the first three modules have shown superior performance in creating high-quality arguments, the debate construction is based on a human-defined template, which cannot adapt to the dynamics in competitive debate. Instead of generating a better argument, our work focuses on the decision-making in the debate, such as which point to argue.
Debate in Large Language Models. There are two research directions in debate with large language models. One is to enhance LLMs’ capability in solving problems by debating with different solutions, such as debate for reasoning (Du et al., 2023; Liang et al., 2023; Yin et al., 2023), evaluation (Chern et al., 2024; Chan et al., 2024), or safety (Irving et al., 2018). Unlike competitive debating, these problems have an optimal solution, and the purpose of the debate is to find this solution by discussion. Another line of work uses LLMs’ argumentation capability to enhance the debating performance. For example, Lee et al. (2023) uses role-playing to simulate debate between different populations. Zhang et al. (2024) design a multi-agent framework to work collaboratively during the debate, showing comparable performance with human debaters. Different from previous work, we consider the strict time limitation in competitive debate, which requires LLMs to make tackle decisions and allocate the time budget to the most important actions.
Language Agents for Games. Besides debating, recent studies also investigate language agents in other strategic games that require strong communication between players, such as Diplomacy (Meta et al., 2022), Werewolf (Xu et al., 2023, 2024), and Avalon (Wang et al., 2023). Cicero (Meta et al., 2022) and (Xu et al., 2024) learn a policy to choose suitable strategies during the game, while Xu et al. (2023) and Wang et al. (2023) design deliberate prompts to improve the performance. These games have a clear definition of winning. For example, the winning condition of villagers in Werewolf is to eliminate all the werewolves, while the winning condition of werewolves is to kill all villagers. These clear winning decisions provide an objective reward signal for learning the strategies. However, there is no clear winner in competitive debate, making the strategy learning more difficult.
@@ 3.1 Competitive Debate
Following previous work on the debate game, we focus on the simplified Oxford-style debate (Slonim et al., 2021; Zhang et al., 2016, 2024). It is a competitive debate format where two sides argue for (Pro stance) and against (Con stance) a specific motion. There are three stages. In Opening Stage, each debater has 4 minutes to deliver their main claims and arguments to support their assigned stance. In Rebuttal Stage, the debaters take turns attacking the opponent and defending themselves in 4 minutes. In Closing Stage, the debaters summarize the debate in 2 minutes.
Audience members are asked to vote before and after listening to the debate. The final winner of the debate is the one with more votes swaying towards it. We call this win as Opinion Shift Win. There are several other debate terms. For example, a claim $c$ is a proposition about the motion, the argument $x$ uses reasoning or evidence to support the claim, and statement is the whole passage presented in one debate stage. There are four typical actions that the debater can take in each stage: propose a new claim, reinforce the existing claims, attack the opponent’s claims, and rebut the opponent’s attack. More detailed descriptions are in Appendix A.
Figure 1: TreeDebater keeps two debate flow trees to track the status for its own side and the opponent’s side. TreeDebater obtains candidate actions from the debate flow tree and retrieves similar nodes with prepared materials from the rehearsal trees. The LLM writer then drafts the statement. TreeDebater also retrieves similar human debate flow trees to help the simulated audience provide revision feedback. The speech time controller is used to control the final speaking time.
@@ 3.2 Rehearsal Tree
Build Rehearsal Tree via Top-Down Specifically, TreeDebater proposes $n$ candidate main claims $C=\{c_{0},\cdots,c_{n}\}$ and constructs a rehearsal tree $T_{r}$ with the maximum depth $L$ for each claim $c=x^{(0)}$. Each node is an argument $x$, and its children are the potential counterarguments. Thus, each node is on the same side of this grandparent. We define the attack score of the $l$-th level node $x^{l}$ as its attack impact towards its parent $x^{l-1}$, which is $r_{a}(x^{l},x^{l-1})$. The support score of $x^{l}$ is its support impact towards its grandparent $x^{l-2}$, which is $r_{s}(x^{l},x^{l-2})$. Here, $r_{s}$ and $r_{a}$ are two scoring models that evaluate the impact between two arguments. For the main claim $x^{0}$ of the root, its support score is defined as its support in the assigned stance $s$.
@@
where $\gamma$ is the decay coefficient since one may choose to stop further reacting, and $\text{Child}(x^{l})$ is the set of $x^{l}$’s child nodes. The whole process of constructing the rehearsal tree is illustrated in Alg. 1.
@@ 3.3 Debate Flow Tree
Maintain Debate Flow Tree The debate flow tree tracks the debate status by keeping all proposed claims with the corresponding attack and defense in a tree structure. The tree node consists of the claim, the arguments to support the claim, the state (proposed or attacked), and the number of visits. Each time after listening to one statement, TreeDebater will update the debate flow tree with the claims and arguments in it. If a new claim occurs, a new node with proposed state will be created. If an existing claim is attacked, the corresponding claim node will change to attacked state, and a child node will be created for this attack. The update process is described in Alg. 2 and Figure 4(b).
Extract Candidate Actions from Debate Flow Tree TreeDebater can filter out the candidate actions it can take in the current stage of the debate based on the debate flow tree. For example, it can rebut the opponent’s attack on its claim by traversing the leaf nodes of the opposite side, or reinforce its existing claims by selecting from the nodes of its side. It can also identify the importance of each action according to the node’s visit number.
Retrieve Prepared Arguments from Rehearsal Tree After getting the candidate actions, TreeDebater retrieves the prepared arguments for this action from the rehearsal trees. For example, to propose or reinforce a claim, we should retrieve the arguments that support this claim or oppose its counterclaim. To attack or rebut a claim, we should retrieve its counterclaim. TreeDebater also retrieves the $k$-step strength score of the claim, and here $k$ depends on the number of rounds remaining in the debate. For example, there are 3 remaining rounds (Con’s opening, Pro’s Rebuttal, Con’s Rebuttal)  [^1] Since the Closing Stage only summarize what have been said, we do not view it as an effective round which can take debate action. for the Pro side in the opening stage, which means that there can be at most 3 levels of child nodes. Therefore, we use $k=3$ to get the strength score. Details of the retrieval process are given in the Appendix B.1.
@@ 3.4 Audience Feedback based on Human Debate Flow Tree
We provide simulated general audience feedback by evaluating different key aspects of the speech based on the retrieved human debate flow trees. These human debate flow trees give the audience a better sense of the back-and-forth and the statement styles in the human debate competition. This helps the simulated audience provide concrete and focused feedback on important aspects in debate, such as message clarity, engagement impact, evidence presentation, and persuasive elements. By revising based on the audience feedback, TreeDebater can learn the allocation pattern and persuasive elements embedded within relevant human debate structures. Details of the construction and retrieval of human debate flow trees are in Appendix D, and an example of the resulting audience feedback is in Table 12.
@@ 3.5 Speech Time Controller
With the selected actions from the debate flow tree, our hope is then to generate a coherent speech within a reasonable time range $[t_{l},t_{r}]$. When drafting the statement, we incorporate the word budget into the instruction, which is based on a rough estimate of the speaking time, i.e., 130 words per minute. However, we find that it is difficult for LLMs to follow the length constraint (Jie et al., 2024; Butcher et al., 2025), and the number of words cannot control the precise speech time since each word takes a different speaking time due to its phoneme and might be affected by emotion. To ensure that the speaking time of the statement fits the precise time limitation in the competitive debate, we introduce a speech time controller in TreeDebater to provide a better estimation of the speech length and guide the LLM writer in revising the statement for the time restriction.
We use a light-weight text-to-speech model, FastSpeech (Ren et al., 2020), to estimate the speech length. At each iteration, the speech time controller converts the current statement to audio and calculates its time cost. The calculated time cost $t$ and the word budget $n$ in the instruction of this statement will be used to search for a new word budget for the revision. The iteration will stop when the speech time falls into the suitable range or it reaches the maximum revision limit.
For the iteration process, we employ a binary search approach to efficiently find an appropriate word count target $n$. Since we observe the real speech time $t$ generally has a positive correlation with $n$, we can find the initial search interval $[n_{l},n_{r}]$ where $n_{l}$ produces a speech shorter than $t_{l}$ and $n_{r}$ produces one longer than $t_{r}$, making the interval dividing process possible. More detailed descriptions of the process are in Appendix B.2.

### v2

@@ Abstract
Winning competitive debates requires sophisticated reasoning and argument skills. There are unique challenges in the competitive debate: (1) The time constraints force debaters to make strategic choices about which points to pursue rather than covering all possible arguments; (2) The persuasiveness of the debate relies on the back-and-forth interaction between arguments, which a single final game status cannot evaluate. To address these challenges, we propose TreeDebater, a novel debate framework that excels in competitive debate. We introduce two tree structures: the Rehearsal Tree and Debate Flow Tree. The Rehearsal Tree anticipates the attack and defenses to evaluate the strength of the claim, while the Debate Flow Tree tracks the debate status to identify the active actions. TreeDebater allocates its time budget among candidate actions and uses the speech time controller and feedback from the simulated audience to revise its statement. The human evaluation on both the stage-level and the debate-level comparison shows that our TreeDebater outperforms the state-of-the-art multi-agent debate system, with a +15.6% improvement in stage-level persuasiveness with DeepSeek and +10% debate-level opinion shift win. Further investigation shows that TreeDebater shows better strategies in limiting time to important debate actions, aligning with the strategies of human debate experts.
@@ 1 Introduction
Competitive debate is a structured battleground of arguments. It plays a crucial role in domains such as education [undefl, undefa], legislation [undefx], and politics [undefi, undefj]. In a debate competition, two sides argue on the same motion under strict time limits. It requires complex reasoning skills and precise selection of impactful arguments. Expert debaters allocate time wisely among arguments in different stages. Unlike other tasks, there is no ground-truth answer in a debate: The winner is determined by who sways the audience or judges more effectively. Research shows that success in debates depends heavily on a debater’s ability of intense back-and-forth competition and on-the-fly decision-making: predict, respond to, and adjust based on their opponent’s arguments [undefy, undefag, undefaa]. Although large language models (LLMs) are effective in generating arguments, human evaluators continue to rate AI debaters as less convincing than human opponents in live debates [undefh]. These gaps highlight that current LLMs struggle with planning effective arguments in the dynamic back-and-forth interaction of competitive debate.
@@
The main challenges of debate AI lie in strict time limits and the lack of objective reward signals. First, the timed setting makes it impossible for LLMs to generate verbose arguments for each candidate action. For example, there can be a trade-off between attacking the opponent’s claims and defending one’s own claims. A formal competitive debate typically consists of three stages: an opening statement, a rebuttal statement, and a closing statement. In each stage, the debater is required to carefully allocate their limited speaking time to propose new claims, reinforce the existing claims, attack the opponent’s claims, or rebut the opponent’s attack. The debaters must prioritize the most impactful actions and dynamically adjust their plan as the debate progresses. However, unlike games such as Go [undefab, undefac] or Werewolf [undefah, undefai], there are no rule-based winning conditions. Instead, the winner depends on the evolving flow of arguments and counterarguments between the two sides. This dynamic nature makes it difficult to plan based on the final reward signals.
@@
In this work, we propose to model the dynamic interaction on trees and show that planning on these trees can significantly improve the debating capacity of LLMs. Human debate experts implicitly employ tree-like reasoning: they rehearse potential rebuttals before the debate starts by anticipating the claims an opponent might say and formulating the potential responses for each of them in a tree format. During the debate, they track the flow of the debate with another tree to keep a structured mental map of which points have been addressed or remain standing. Inspired by these human strategies, we introduce a new debate system TreeDebater to help LLMs make tactical decisions in the debate. Akin to human debaters, we introduce two kinds of trees in TreeDebater: Rehearsal Tree to hypothesize the opponent’s attacks in advance, and Debate Flow Tree to track the debate status. Specifically, in the Rehearsal Trees, TreeDebater scores arguments based on the potential tree-structured attacks and defenses. During the debate, TreeDebater keeps track of the debate status through the Debate Flow Trees and provides the candidate action set. TreeDebater retrieves prepared arguments from the Rehearsal Tree for each candidate action to draft the statement. Finally, TreeDebater revises its statement based on simulated audience feedback and uses a speech-based search algorithm to fit the statement into the speaking time limitation.
@@
To evaluate the effectiveness of TreeDebater in competitive debate, we carried out extensive experiments with the multi-agent debate system Agent4Debate [undefal], which has shown comparable results to human debaters. We design two kinds of human evaluation to compare the debate performance: stage-level head-to-head comparison with a fixed debate context, and debate-level end-to-end comparison with an Oxford-style debate format. Human evaluation demonstrates that TreeDebater consistently outperforms the baseline in the dimension of the average persuasiveness score and the stage-level and debate-level win rate. It is preferred 1.5x and 3.5x over the baseline in the stage-level and debate-level comparison with the Gemini-2.0-flash backbone, where the gain also generalizes to another backbone, DeepSeek-V3.
@@ 2 Related Work
Computational Argumentation. Previous studies used computational methods to study argumentation processes, such as argument mining [undefo], the polarity of arguments [undef, undefv, undefam], argument structure [undefak, undefq], and argument generation [undefk, undeft]. [undefad] built the first autonomous debating system, Project Debater, with four argumentation modules: argument mining, argument knowledge base, argument rebuttal, and debate construction. While the first three modules have shown superior performance in creating high-quality arguments, the debate construction is based on a human-defined template, which cannot adapt to the dynamics in competitive debate. Instead of generating a better argument, our work focuses on the decision-making in the debate, such as which point to argue.
Debate in Large Language Models. There are two research directions in debate with large language models. One is to enhance LLMs’ capability in solving problems by debating with different solutions, such as debate for reasoning [undefe, undefs, undefaj], evaluation [undefd, undefc], or safety [undefm]. Unlike competitive debating, these problems have an optimal solution, and the purpose of the debate is to find this solution by discussion. Another line of work uses LLMs’ argumentation capability to enhance the debating performance. For example, [undefp] uses role-playing to simulate debate between different populations. [undefal] design a multi-agent framework to work collaboratively during the debate, showing comparable performance with human debaters. Different from previous work, we consider the strict time limitation in competitive debate, which requires LLMs to make tackle decisions and allocate the time budget to the most important actions.
Language Agents for Games. Besides debating, recent studies also investigate language agents in other strategic games that require strong communication between players, such as Diplomacy [undefw], Werewolf [undefah, undefai], and Avalon [undefaf]. Cicero [undefw] and [undefai] learn a policy to choose suitable strategies during the game, while [undefah] and [undefaf] design deliberate prompts to improve the performance. These games have a clear definition of winning. For example, the winning condition of villagers in Werewolf is to eliminate all the werewolves, while the winning condition of werewolves is to kill all villagers. These clear winning decisions provide an objective reward signal for learning the strategies. However, there is no clear winner in competitive debate, making the strategy learning more difficult.
@@ 3.1 Competitive Debate
Following previous work on the debate game, we focus on the simplified Oxford-style debate [undefad, undefak, undefal]. It is a competitive debate format where two sides argue for (Pro stance) and against (Con stance) a specific motion. There are three stages. In Opening Stage, each debater has 4 minutes to deliver their main claims and arguments to support their assigned stance. In Rebuttal Stage, the debaters take turns attacking the opponent and defending themselves in 4 minutes. In Closing Stage, the debaters summarize the debate in 2 minutes.
Audience members are asked to vote before and after listening to the debate. The final winner of the debate is the one with more votes swaying towards it. We call this win as Opinion Shift Win. There are several other debate terms. For example, a claim $c$ is a proposition about the motion, the argument $x$ uses reasoning or evidence to support the claim, and statement is the whole passage presented in one debate stage. There are four typical actions that the debater can take in each stage: propose a new claim, reinforce the existing claims, attack the opponent’s claims, and rebut the opponent’s attack. A battlefield refers to a conflict zone in the debate, where two sides contest on a specific point. It includes several rounds of proposal, attack, and defense.
Figure 1: The overall workflow. In each stage, TreeDebater (i) updates two Debate Flow Trees with extracted action tuples from the current statement; (ii) retrieves prepared arguments from the Rehearsal Tree for the candidate actions; (iii) generates a draft based on the retrieved arguments and important scores; (vi) lets simulated audience provide feedback based on retrieved human debate flow tree; (v) revises based on feedback from the simulated audience and speech time controller.
@@ 3.2 Rehearsal Tree
Build Rehearsal Tree via Top-Down Specifically, TreeDebater proposes $n$ candidate main claims $C=\{c_{0},\cdots,c_{n}\}$ and constructs a Rehearsal Tree $T_{r}$ with the maximum depth $L$ for each claim $c=x^{(0)}$. Each node is an argument $x$, and its children are the potential counterarguments. Thus, each node is on the same side of its grandparent. We define the attack score of the $l$-th level node $x^{l}$ as its attack impact towards its parent $x^{l-1}$, which is $r_{a}(x^{l},x^{l-1})$. The support score of $x^{l}$ is its support impact towards its grandparent $x^{l-2}$, which is $r_{s}(x^{l},x^{l-2})$. Here, $r_{s}$ and $r_{a}$ are two scoring models that evaluate the impact between two arguments. For the main claim $x^{0}$ of the root, its support score is defined as its support in the assigned stance $s$.
@@
where $\gamma$ is the decay coefficient (which is 0.8) since one may choose to stop further reacting, and $\text{Child}(x^{l})$ is the set of $x^{l}$’s child nodes. The whole process of constructing the Rehearsal Tree is illustrated in Alg. 1.
@@ 3.3 Debate Flow Tree
Maintain Debate Flow Tree The Debate Flow Tree tracks the debate status by keeping all proposed claims with the corresponding attack and defense in a tree structure. The tree node consists of the claim, the arguments to support the claim, the state (proposed or attacked), and the number of visits. Each time after listening to one statement, TreeDebater first extracts the tuples of (action, claim, argument, target) from the statement, and then updates the Debate Flow Tree with these tuples. If a new claim occurs, a new node with proposed state will be created under the tree root.. If an existing claim is attacked, its target claim node will change to attacked state, and a child node will be created to indicate this attack. When new arguments are created to reinforce an existing claim, the corresponding claim node will be updated with these new arguments. The update process is described in Alg. 2 and Figure 4(b).
Extract Candidate Actions from Debate Flow Tree TreeDebater can filter out the candidate actions it can take in the current stage of the debate based on the Debate Flow Tree. The criteria include: (a) propose only at the opening stage; (b) rebut the latest nodes (leaf node) of the opposite side; (c) reinforce the nodes of its own side; (d) attack the nodes of the opposite side.
Retrieve Prepared Arguments from Rehearsal Tree TreeDebater retrieves the arguments prepared for this action from the Rehearsal Trees after getting the candidate actions. For example, to propose or reinforce a claim, we should retrieve the arguments that support this claim or oppose its counterclaim. To attack or rebut a claim, we should retrieve its counterclaim. TreeDebater also retrieves the $k$-step strength score of the claim, and here $k$ depends on the number of rounds remaining in the debate. For example, there are 3 remaining rounds (Con’s opening, Pro’s Rebuttal, Con’s Rebuttal)  [^1] Since the Closing Stage only summarize what have been said, we do not view it as an effective round which can take debate action. for the Pro side in the opening stage, which means that there can be at most 3 levels of child nodes. Therefore, we use $k=3$ to get the strength score. We use one of the state-of-the-art embedding models, Gemini-text-embedding-4 to get the embedding of the claims. If the cosine similarity of two embeddings is higher than a threshold of 0.8, we view these two claims are similar. To find the most similar tree node, we traverse all tree nodes and calculate the cosine similarity between the claim of the node and the target claim. Details of the retrieval process are given in Appendix D.
@@ 3.4 Audience Feedback based on Human Debate Flow Tree
We provide simulated general audience feedback by evaluating different key aspects of the speech based on the retrieved human Debate Flow Trees. We first curate a human debate flow tree corpus with Debate Flow Tree created from human debate corpus. During the debate, we reformulate the current Debate Flow Tree into a tree-like string. Then we employ semantic search based on the embeddings of the debate flow tree string. Similarly to the argument retrieval, Gemini-text-embedding-4 is used to get the embedding of the tree-like string of the current debate flow tree and of trees in human debate flow tree corpus. The top 1 human debate flow tree is retrieved with a similar threshold of 0.8. We add the retrieval to the instruction for simulated audience to give the audience a better sense of the back-and-forth and the statement styles in the human debate competition.
The simulated audience provide concrete and focused feedback on important aspects in debate, such as message clarity, engagement impact, evidence presentation, and persuasive elements. By revising based on the audience feedback, TreeDebater can learn the allocation pattern and persuasive elements embedded within relevant human debate structures. Details of the construction and retrieval of human Debate Flow Trees are in Appendix D, and an example of the resulting audience feedback is in Table 15.
@@ 3.5 Speech Time Controller
With the selected actions from the Debate Flow Tree, our hope is then to generate a coherent speech within a reasonable time range $t_{l},t_{r}]$. When drafting the statement, we incorporate the word budget into the instruction, which is based on a rough estimate of the speaking time, i.e., 130 words per minute. However, we find that it is difficult for LLMs to follow the length constraint [undefn, undefb], and the number of words cannot control the precise speech time since each word takes a different speaking time due to its phoneme and might be affected by emotion. To ensure that the speaking time of the statement fits the precise time limitation in the competitive debate, we introduce a speech time controller in TreeDebater to provide a better estimation of the speech length and guide the LLM writer in revising the statement for the time restriction.
We use a light-weight text-to-speech model, FastSpeech [undefz], to estimate the speech length. At each iteration, the speech time controller converts the current statement to audio and calculates its time cost. The calculated time cost $t$ and the word budget $n$ in the instruction of this statement will be used to search for a new word budget for the revision. The iteration will stop when the speech time falls into the suitable range or it reaches the maximum revision limit.
For the iteration process, we employ a binary search approach to efficiently find an appropriate word count target $n$. Since we observe the real speech time $t$ generally has a positive correlation with $n$, we can find the initial search interval $n_{l},n_{r}]$ where $n_{l}$ produces a speech shorter than $t_{l}$ and $n_{r}$ produces one longer than $t_{r}$, making the interval dividing process possible. More detailed descriptions of the process are in Appendix B.2.